OpenStudy (insa):

in an experiment, a small spherical drop of oil is allowed to fall onto the surface of water so that it produces a thin film of oil covering a large area. a) given that the volume of a drop of oil is 12.5mm^3, find the number of drops which can be produced by 5000mm^3 of oil. b) given that the volume V of a sphereof radius 4/3(pie)r^3, express r in terms of V and (pie). then calculate the radius of one drop of oil.

OpenStudy (insa):

@paki need help again!

OpenStudy (anonymous):

haha tummhai toh help e chahiye hoti hai khud b kuch kr liya karo nalaaik..... :P

OpenStudy (paki):

welcome @insa i am here now.... :-) office se abhi chutti mili hai.... :-)

OpenStudy (paki):

part A of your question is very easy..... just look it... ok.... here the volume of a drop of oil is 12.5 mm^3... so the number of drops which can be produced by 5000 mm^3 of oil will be ===> 5000/12.5 = 400 drops....

OpenStudy (paki):

and for part B... we will use some formulas to calculate the required term.... see... here we have the volume V of a sphere of radius r is 4/3 pi r^3, expressed r in terms of V and pi. Then calculate the radius of the drop of oil from the relation..... \[v = 4 * \pi * r ^ {3} \div 3\]from here take out "r".... which is.... \[r ^{3} = 3 *v \div 4 * \]now remove the cube from "r", so that the equation may be simple.... \[r = (3 * v \div 4* \pi) ^{1/3}\] now this is our equation... just put the values... \[r = ( 3 * 12.5 \div 4 * \pi) ^{1/3}\]we will get "r" now... r= 1.43 mm

OpenStudy (insa):

@roha tm to meray se text mein baat kro na,, btati hn tmhay...

OpenStudy (insa):

thx alot again :) @paki

OpenStudy (anonymous):

hahaha

OpenStudy (paki):

@insa pleasure.... :-)